-7/3u+4/5=-2/3u-4
15 (-7/3u+4/5)=15 (-2/3u-4)
15 (-7/3u)+15 (4/5)=2/3u)-15 (4)
-35u+12=-10u-60
-35u+10u+12=60
-35u+10u=-60-12
-45u=-72
U=-72/-455
U=-5/-8
Please help with question above the working I get correct until it'stime to separate left from right side. to add or subtract? When should the numbers be positive or negative?
/ denotes -7 over
Thank you

Found 2 solutions by Cromlix, Alan3354:
Answer by Cromlix(4381)

Hi there,
-7/3u+4/5=-2/3u-4
Collect like terms
-7/3u + 2/3u = -4 - 4/5
-5/3u = -20/5(4) - 4/5
-5/3u = -24/5
Cross multiply
-5 x 5 = 3u x -24
-25 = -72u
u = 25/72
Hope this helps:-)

Answer by Alan3354(69443)

-7/3u+4/5=-2/3u-4
15 (-7/3u+4/5)=15 (-2/3u-4)
15 (-7/3u)+15 (4/5)=2/3u)-15 (4)
-35u+12=-10u-60
-35u+10u+12=60
-35u+10u=-60-12
-45u=-72 **************** -25u = -72
U=-72/-455 ????? 455
U=-5/-8
Please help with question above the working I get correct until it'stime to separate left from right side. to add or subtract? When should the numbers be positive or negative?
/ denotes -7 over
===================
It appears you mean
(-7/3)u + 4/5 = (-2/3)u - 4
Multiply by 15
-35u + 12 = -10u - 60
72 = 25u
u = 72/25